Grade 12th passMechanics

Displacement, x (in meters), of a body of mass 1 kg as a function of time, t, on a horizontal smooth surface is give as x = 2t2 . The work done in the first one second by the external force is

Abin Santhosh
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
x = 2 t^2
 
dx / dt = v =  4t
 
d2x / dt2  = a =  4
 
Force acting on a body = m*a = 1 * 4 = 4 Newton
 
 W = \int F . dx
 
w = \int 4 * 4t dt
 
 = 16 t^2 / 2 = 8 t^2
 
now put t = 1
 
hence W = 8 Joules